# MIND-V: Hierarchical Video Generation for Long-Horizon Robotic Manipulation with RL-based Physical Alignment

### v1

@@ 3.2 Test-Time Optimization via Staged Visual Future Rollouts
Long-horizon task generation is plagued by error accumulation, where minor sub-task deviations in one sub-task cascade into overall task failure [robridge]. To mitigate this risk, we introduce a novel test-time optimization strategy: Staged Visual Future Rollouts. This strategy operates through a dynamic cycle of proposal, verification, and refinement executed at each sub-task transition. This approach effectively decomposes the global planning challenge into a series of locally optimal decisions, thereby preventing catastrophic error propagation.
As illustrated in Figure 2(d), the process begins with the SRH proposing a set of $K$ semantically plausible yet strategically diverse candidate trajectories. These trajectories are then synthesized by the MVG into corresponding video clips, $V_{K}$, each depicting a potential future outcome. Subsequently, the VLM transitions into a verification judge, evaluating each candidate future based on criteria including task success, physical plausibility, and visual quality.
If the highest-scoring video $V_{top}$ meets a predefined success threshold, it is selected and the process proceeds. If no candidate is satisfactory, the VLM provides structured textual feedback detailing the failure modes (e.g., “end position error”, “object not grasped correctly”). This feedback loop instructs the SRH to re-plan and propose a refined set of masks and trajectories in the next iteration. This iterative cycle of propose-verify-refine transforms the SRH from a simple feed-forward planner into a proactive, self-correcting agent, significantly enhancing the robustness and success rate of long-horizon task generation.
Figure 4: Physical Foresight Coherence (PFC) Reward. The PFC leverages a frozen V-JEPA2 world model to predict the latent representation of future Target frames conditioned on past Context frames. The reward is the cosine similarity between this prediction and the ground-truth target latent, which measures the video’s alignment with the world model’s learned physical dynamics.
3.3 MVG Training: From Supervised Fine-Tuning to Physical Alignment
To ensure the MVG generates high-fidelity videos, we employ a two-stage training paradigm. First, Supervised Fine-Tuning (SFT) adapts a pre-trained video model to the robotics domain. This is followed by a GRPO Reinforcement Learning (RL) post-training phase that aligns the generator with abstract objectives such as physical plausibility and aesthetic quality.
3.3.1 Supervised Fine-Tuning (SFT)
The SFT phase adapts an open-source video model to learn the fundamental mapping from our structured Behavioral Semantic Bridge (BSB) representation to coherent video sequences. We fine-tune the model on a real-world robotics dataset (e.g., Bridge v2 [bridge]) using ground-truth BSB annotations. The optimization follows the standard denoising objective, conditioned on the BSB:
@@ 3.3.1 Supervised Fine-Tuning (SFT)
where $x_{t}$ is the noised video at timestep $t$. This stage provides a high-quality initial policy $\pi_{\text{ref}}$ for the subsequent alignment phase. Notably, the model only needs to be trained on short sub-task videos, as the hierarchical framework enables generalization to arbitrary long-horizon tasks.
3.3.2 GRPO Post-Training
While SFT enforces motion adherence, it cannot guarantee physical plausibility or aesthetic quality of the generated videos, objectives that are ill-suited for conventional loss functions [flow_grpo]. To bridge this gap, we introduce a post-training alignment stage using Reinforcement Learning (RL). This stage models the denoising process as a Markov Decision Process (MDP) and employs Group Relative Policy Optimization (GRPO) [flow_grpo] for optimization.
The optimization is guided by a composite reward function $R(x_{0})$, which evaluates the overall quality of a generated video $x_{0}$ by taking a weighted sum of a physics-based reward and an aesthetic reward:
$R(x_{0})=w_{p}\cdot R_{\text{physics}}(x_{0})+w_{a}\cdot R_{\text{aesthetic}}(x_{0}).$ (3)
Physical Foresight Coherence (PFC) Reward ($R_{\text{physics}}$): This reward innovatively employs a pre-trained visual world model, V-JEPA2 [vjepa2], to serve as an objective arbiter of physical plausibility (Figure 4). Pre-trained via self-supervision on large-scale real-world data and fine-tuned on a robotics dataset, V-JEPA2 acquires an internal model of world dynamics, enabling it to accurately understand the existing state and predict future states within an abstract latent space. For each generated video, a sliding window approach is employed to assess local physical plausibility. The consistency score $s_{i}$ for each window is the cosine similarity between V-JEPA2’s latent prediction and the actual future:
@@ 3.3.2 GRPO Post-Training
Aesthetic Reward ($R_{\text{aesthetic}}$): The aesthetic reward is provided by a VLM, such as Qwen-VL [Qwen-VL], which performs tiered scoring. The VLM assesses each video for clarity, artifacts, and realism, assigning a discrete integer score (e.g., 1-5), which yields a stable and discriminative reward signal for optimization.
GRPO Optimization GRPO is an efficient, value-free policy optimization algorithm. At each optimization step, we sample a group of $G$ videos $\{x_{0}^{i}\}_{i=1}^{G}$ from the current policy $\pi_{\theta}$. The advantage $\hat{A}^{i}$ for each sample is then computed by normalizing its reward relative to the group’s statistics:
@@
where $\text{mean}(\cdot)$ and $\text{std}(\cdot)$ are the mean and standard deviation of the rewards within the group. The policy is then updated by maximizing the GRPO objective function:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\bigg($ $\min\left(r_{i}(\theta)\hat{A}^{i},\text{clip}(r_{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}^{i}\right)$
$\hskip 18.49988pt-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg)\Bigg],$ (7)
where $r_{i}(\theta)=\frac{\pi_{\theta}(x_{0}^{i})}{\pi_{ref}(x_{0}^{i})}$ is the importance sampling ratio, $\epsilon$ is a clipping hyperparameter, and the KL-divergence term regularizes the policy towards the SFT policy $\pi_{\text{ref}}$ to mitigate reward hacking. This optimization process progressively aligns the generator towards higher physical fidelity and aesthetic quality while maintaining its adherence to kinematic conditioning.
@@
4 Experiments
Table 1: Visual quality evaluation on short-horizon and long-horizon tasks. Our model is benchmarked against a consistent set of state-of-the-art methods across both task types. Higher values are better and highlighted in shades of green.
Method Aesthetic Quality  $\uparrow$ Imaging Quality  $\uparrow$ Temporal Flickering  $\uparrow$ Motion Smoothness  $\uparrow$ Subject Consistency  $\uparrow$ Background Consistency  $\uparrow$
Short-horizon Task Evaluation
@@ 4 Experiments
Long-horizon Task Evaluation
@@
Wan2.2-14B [wan22] 0.508 0.661 0.948 0.951 0.885 0.913
HunyuanVideo [hunyuanvideo] 0.487 0.643 0.928 0.954 0.862 0.900
MIND-V (Ours) 0.504 0.658 0.955 0.953 0.896 0.924
Table 2: Comprehensive evaluation of long-horizon tasks on PFC Score, Task Completion Rates, and User Preferences. Higher values are better and highlighted in shades of green.
Method
PFC Score
$\uparrow$
Task Success
Rate (%) $\uparrow$
User Study
(Preference %) $\uparrow$
Robodreamer [robodreamer] 0.418 27.5 6.7
WoW-1-DiT-7B [wow] 0.423 32.2 16.7
WoW-1-Wan-14B [wow] 0.420 34.7 23.3
Wan2.2-14B [wan22] 0.402 11.1 0
HunyuanVideo [hunyuanvideo] 0.411 9.8 3.3
MIND-V (Ours) 0.445 61.3 46.7
4.1 Experiment Settings
Architecture and Training  Our Semantic Reasoning Hub (SRH) employs Gemini-2.5 Pro API [gemini25] as its core Vision-Language Model (VLM), complemented by Affordance-R1 [affordance-r1] as visual localizer. The Motor Video Generator (MVG) is initialized from the pre-trained CogVideoX-5B [cogvideox] architecture. Experiments are conducted on the Bridge V2 [bridge] dataset, following the data processing protocol established in [robomaster]. We adopt a resolution of $480\times 640$ pixels and a video length of 37 frames per sub-task for both training and inference. The model underwent two training stages: (1) Supervised fine-tuning (SFT) for 30,000 steps with an AdamW optimizer and a learning rate of $2\times 10^{-5}$; and (2) GRPO post-training for 1,500 iterations at a learning rate of $5\times 10^{-5}$. At inference time, MIND-V can generate a 111-frame long-horizon video (comprising three sub-tasks) in approximately 180 seconds while consuming around 50 GB of VRAM. Theoretically, due to its autoregressive, sub-task-based architecture, MIND-V can generate arbitrarily long task-sequence videos with only a linear increase in computational cost. All experiments are conducted on four NVIDIA H200 GPUs. Additional implementation details are provided in the supplementary material.
Evaluation Protocol and Metrics  The evaluation is performed on a test set of 108 samples. This set comprises scenes from the Bridge V2 test set [bridge] and unseen scenes sourced from the web. Recognizing that different task horizons demand different evaluation criteria, we adopted a bifurcated protocol. For short-horizon tasks, our evaluation focuses exclusively on visual quality, where we use V-Bench [vbench2] for evaluation. For long-horizon tasks, we introduce user study and two additional metrics: (1) physical plausibility, which is quantified by our Physical Foresight Coherence (PFC) score, as detailed in Section 3.3.2; and (2) Task Success Rate, which measures the average success rate across all sub-tasks of the entire long-horizon task.
Baselines and Comparative Setup  For short-horizon task, we benchmark MIND-V against both trajectory-based methods (IRASim [IRASim], MotionCtrl [motionctrl], DragAnything [draganything], Tora [tora]) and trajectory-free world models (Robodreamer [robodreamer], WoW [wow], Wan2.2 [wan22], HunyuanVideo [hunyuanvideo]). It is crucial to note that trajectory-based models receive privileged information at inference time, such as manual trajectories, masks, or anchor points, which is unavailable to our trajectory-free approach. This distinction underscores the greater complexity of the task our model addresses. For long-horizon task, which emphasizes complex planning and reasoning without explicit guidance, our comparison is focused on SOTA trajectory-free models (Robodreamer [robodreamer], WoW [wow] and WAN2.2 [wan22], HunyuanVideo [hunyuanvideo]). Each long-horizon task in our benchmark is composed of a sequence of 2 to 4 sub-tasks, designed to probe the limits of long-horizon planning and generation.
@@ 4.2 Qualitative and Quantitative Comparison
As evidenced by the results in Figure 5, Table 1, and Table 2, MIND-V consistently outperforms state-of-the-art methods across both short- and long-horizon tasks. While baseline models may achieve high scores on visual quality metrics, they consistently struggle with task execution and long-horizon coherence (Table 2). These models exhibit critical deficiencies (Figure 5), including logical failures, such as hallucinating unprompted actions; physical implausibility, like objects spontaneously disappearing; and poor semantic grounding, which leads to inaction or incorrect object manipulation.
Figure 6: Visualization of the SRH Planning.
In contrast, MIND-V successfully executes these long-horizon tasks, demonstrating robust adherence to both instructions and physical principles. This superior performance is attributed to our model’s hierarchical design: The Semantic Reasoning Hub (SRH) and Behavioral Semantic Bridge (BSB) collaborate to decompose user instructions into an explicit, executable plan (Figure 6), which mitigates the risk of semantic drift common in end-to-end models. Subsequently, the Motor Video Generator (MVG), guided by this plan and fine-tuned with our Physical Foresight Coherence (PFC) reward, ensures the resulting synthesis adheres to physical laws. This systematic framework of cognition from execution effectively prevents the error accumulation that plagues other methods, enabling the generation of coherent and physically plausible manipulation sequences.

### v2

@@ 3.2 Test-Time Optimization via Staged Visual Future Rollouts
Simulating long-horizon futures is plagued by error accumulation, where minor subtask deviations in one subtask cascade into overall task failure [robridge]. To mitigate this risk, we reframe inference as a search process in the world model’s latent space. We introduce Staged Visual Future Rollouts, a novel test-time optimization strategy that mimics “System 2” thinking [system2] by deliberating over potential futures at critical decision points.
As illustrated in Fig. 2(d), this mechanism operates as a “Propose-Verify-Refine” loop at each subtask transition. Instead of blindly autoregressing pixels, the SRH first proposes $K$ semantically plausible candidate plans (i.e., diverse BSBs). The MVG then acts as a parallel simulator, rolling out these plans into short-horizon video clips $\{V_{1},\dots,V_{K}\}$. Crucially, the VLM transitions from a planner to a Critic, evaluating each rollout based on task success, physical plausibility, and visual quality.
If the highest-scoring video $V_{top}$ meets a predefined success threshold, it is selected and the process proceeds. If no candidate is satisfactory, the VLM provides structured textual feedback detailing the failure modes (e.g., “end position error”, “object not grasped correctly”). This feedback loop instructs the SRH to re-plan and propose a refined set of masks and trajectories in the next iteration. This iterative cycle of propose-verify-refine transforms the VWM from a simple feed-forward predictor into a proactive, self-correcting agent, significantly enhancing the robustness and success rate of long-horizon task simulations.
Figure 4: (a) Illustration of the SRH Planning. (b) Physical Foresight Coherence (PFC) Reward. The PFC leverages a frozen V-JEPA2 world model as a “Physics Referee” to predict the latent evolution of future states. The reward quantifies the alignment between the generated video’s dynamics and V-JEPA2’s internal physical laws, computed via cosine similarity in the predictive latent space.
3.3 MVG Training: From SFT to RL Alignment
To establish MVG as a physically grounded simulator, we employ a two-stage training paradigm. First, Supervised Fine-Tuning (SFT) adapts a pre-trained video model to the robotics domain. This is followed by a GRPO Reinforcement Learning (RL) post-training phase that aligns the model’s internal dynamics with fundamental physical laws and aesthetic standards.
Stage 1: Supervised Fine-Tuning (SFT)  The SFT phase adapts a video foundation model to learn the fundamental mapping from the structured Behavioral Semantic Bridge (BSB) to coherent pixel-level execution. We fine-tune the video foundation model on a real-world robotics dataset (e.g., Bridge v2 [bridge]) using ground-truth BSB annotations. The optimization minimizes the standard denoising objective conditioned on the BSB:
@@ 3.3 MVG Training: From SFT to RL Alignment
where $x_{t}$ is the noised video at timestep $t$. This stage provides a high-quality initial policy $\pi_{\text{ref}}$ for the subsequent alignment phase. Notably, the model only needs to be trained on short subtask videos, as the hierarchical framework enables generalization to arbitrary long-horizon tasks.
Stage 2: Physical Alignment via GRPO Post-Training   While SFT enforces motion adherence, it cannot guarantee physical plausibility or aesthetic quality of the generated videos, objectives that are ill-suited for conventional loss functions [flow_grpo]. To bridge this gap, we introduce a post-training alignment stage using Reinforcement Learning (RL). This stage models the denoising process as a Markov Decision Process (MDP) and employs Group Relative Policy Optimization (GRPO) [flow_grpo] for optimization.
The optimization is guided by a composite reward function $R(x_{0})$ that assesses the overall quality of the generated video $x_{0}$ via a weighted combination of a physics-based reward and an aesthetic reward:
$R(x_{0})=w_{p}\cdot R_{\text{physics}}(x_{0})+w_{a}\cdot R_{\text{aesthetic}}(x_{0}),$ (3)
with $w_{p}=0.2$ and $w_{a}=1$. Ablation details are presented in Sec. 4.4.
Physical Foresight Coherence (PFC) Reward ($R_{\text{physics}}$): This reward innovatively employs a powerful pre-trained world model, V-JEPA2 [vjepa2], as a “Physics Referee”. Pre-trained via self-supervision on large-scale real-world data and fine-tuned on robotics datasets, V-JEPA2 acquires an internal model of world dynamics, enabling it to accurately understand the existing state and predict future states within an abstract latent space. For each generated video, a sliding window approach is employed to assess local physical plausibility. The consistency score $s_{i}$ for each window is the cosine similarity between V-JEPA2’s latent prediction and the actual future (Fig. 4):
@@
Aesthetic Reward ($R_{\text{aesthetic}}$): The aesthetic reward is provided by a VLM, such as Gemini 2.5 [gemini25]. The VLM assesses each video for clarity, artifacts, and realism, assigning a discrete integer score (e.g., 1-5), which yields a stable and discriminative reward signal for optimization.
GRPO Optimization: GRPO is an efficient, value-free policy optimization algorithm. At each optimization step, we sample a group of $G$ videos $\{x_{0}^{i}\}_{i=1}^{G}$ from the current policy $\pi_{\theta}$. The advantage $\hat{A}^{i}$ for each sample is then computed by normalizing its reward relative to the group’s statistics:
@@
where $\text{mean}(\cdot)$ and $\text{std}(\cdot)$ denote the mean and standard deviation of the rewards within the group. This group-relative normalization provides a stable advantage estimation without requiring a separate critic network. The policy is then updated by maximizing the standard GRPO objective function:
$\small\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\bigg(\min\left(r_{i}(\theta)\hat{A}^{i},\text{clip}(r_{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}^{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg)\Bigg]$ (7)
where $r_{i}(\theta)=\frac{\pi_{\theta}(x_{0}^{i})}{\pi_{ref}(x_{0}^{i})}$ is the importance sampling ratio, $\epsilon$ is a clipping hyperparameter that ensures conservative policy updates, and the KL-divergence term regularizes the policy towards the SFT policy $\pi_{\text{ref}}$ to mitigate reward hacking and maintain generation quality. This optimization process progressively aligns the MVG towards higher physical fidelity and aesthetic quality while maintaining its adherence to kinematic conditioning.
4 Experiments
4.1 Experiment Settings
Architecture and Training  Our Semantic Reasoning Hub (SRH) employs Gemini-2.5 Pro API [gemini25] as its core Vision-Language Model (VLM), complemented by Affordance-R1 [affordance-r1] as visual localizer. The Motor Video Generator (MVG) is initialized from the pre-trained CogVideoX-5B [cogvideox] architecture. Experiments are conducted on the Bridge V2 [bridge] dataset, following the data processing protocol established in [robomaster]. We adopt a resolution of $480\times 640$ pixels and a video length of 37 frames per subtask for both training and inference. The model underwent two training stages: (1) Supervised fine-tuning (SFT) for 30,000 steps with an AdamW optimizer and a learning rate of $2\times 10^{-5}$; and (2) GRPO post-training for 1,500 iterations at a learning rate of $5\times 10^{-5}$. At inference time, MIND-V can generate a 111-frame long-horizon video (comprising three subtasks) with around 50 GB of VRAM. Theoretically, due to its autoregressive, subtask-based architecture, MIND-V can generate arbitrarily long task-sequence videos with only a linear increase in computational cost. All experiments are conducted on four NVIDIA H200 GPUs. Additional implementation details are provided in the supplementary material.
@@ 4.1 Experiment Settings
Evaluation Protocol and Metrics  The evaluation is performed on a test set of 108 samples. This set comprises scenes from the Bridge V2 test set [bridge] and unseen scenes sourced from the web. Recognizing that different task horizons demand different evaluation criteria, we adopt a bifurcated protocol. For short-horizon tasks, our evaluation focuses exclusively on visual quality, where we use V-Bench [vbench2] for evaluation. For long-horizon tasks, we introduce user study and two additional metrics: (1) physical plausibility, which is quantified by our Physical Foresight Coherence (PFC) score, as detailed in Section 3.3; and (2) Task Success Rate, which measures the average success rate across all subtasks of the entire long-horizon task.
Table 1: Visual quality evaluation on short-horizon and long-horizon tasks. Higher is better; best and second-best results are highlighted.
Method Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$ Temporal Flicker $\uparrow$ Motion Smoothness $\uparrow$ Subject Consistency $\uparrow$ Bg. Consistency $\uparrow$
Short-horizon Tasks
@@
Long-horizon Tasks
@@
Dreamdojo [dreamdojo] 0.504 0.660 0.944 0.948 0.883 0.909
HunyuanVideo [hunyuanvideo] 0.487 0.643 0.928 0.952 0.862 0.900
CogVideoX-5B [cogvideox] 0.484 0.640 0.926 0.947 0.860 0.895
MIND-V (Ours) 0.512 0.658 0.955 0.953 0.896 0.924
Baselines and Comparative Setup  For short-horizon task, we benchmark MIND-V against both autonomous world models (Dreamdojo [dreamdojo], Robodreamer [robodreamer], WoW [wow], CogVideoX-5B [cogvideox], HunyuanVideo [hunyuanvideo]) and trajectory-guided methods (IRASim [IRASim], Cosmos [cosmos], MotionCtrl [motionctrl], DragAnything [draganything], Tora [tora]). Among them, video foundation models such as CogVideoX-5B and HunyuanVideo are fine-tuned on our dataset for fair comparison. It is crucial to note that trajectory-based models receive privileged information at inference time, such as manual trajectories, masks, or anchor points, which is unavailable to our trajectory-free approach. This distinction highlights MIND-V’s capability to infer complex dynamics solely from high-level intent. For long-horizon task, which emphasizes complex planning and reasoning without explicit guidance, our comparison is focused on SOTA autonomous world models (Dreamdojo, Robodreamer, WoW, CogVideoX-5B and HunyuanVideo). Each long-horizon task in our benchmark is composed of a sequence of 2 to 4 subtasks, designed to probe the limits of long-horizon planning and generation.
@@ 4.2 Qualitative and Quantitative Comparison
As demonstrated in Fig. 5 and Tables 12, MIND-V consistently outperforms state-of-the-art baselines, establishing a new standard for both short- and long-horizon robotic simulation tasks. In long-horizon tasks, MIND-V surpasses the second-best method by 9.0%, 76.7%, and 172.2% on PFC Score, Task Success Rate, and User Preference, respectively. Baseline models consistently struggle with task execution and long-horizon coherence (Table 2). They exhibit critical deficiencies (Fig. 5(a)), including logical failures (e.g., hallucinating unprompted actions), physical implausibility (e.g., spontaneous object disappearance), and inaccurate semantic grounding (e.g., incorrect object manipulation).
Table 2: Comprehensive evaluation of long-horizon tasks (PFC Score, Task Success Rate, and User Preferences). Higher is better; best results are bolded.
Method PFC Score $\uparrow$ Task Success Rate User Pref. (%) $\uparrow$
Dreamdojo [dreamdojo] 0.424 0.333 18
Robodreamer [robodreamer] 0.418 0.275 7
WoW-1-DiT-7B [wow] 0.423 0.322 11
WoW-1-Wan-14B [wow] 0.420 0.347 14
CogVideoX-5B [cogvideox] 0.406 0.081 0
HunyuanVideo [hunyuanvideo] 0.411 0.098 1
MIND-V (Ours) 0.462 0.613 49
In contrast, MIND-V successfully executes these long-horizon tasks, demonstrating robust adherence to both instructions and physical principles. This superior performance is attributed to our cognitive hierarchical design: The Semantic Reasoning Hub (SRH) and Behavioral Semantic Bridge (BSB) collaborate to decompose user instructions into an explicit, executable plan, which mitigates the risk of semantic drift common in end-to-end models. Subsequently, the Motor Video Generator (MVG), aligned via with our Physical Foresight Coherence (PFC) reward, synthesizes simulation videos that strictly adhere to physical laws and command requirements. Furthermore, the Staged Visual Future Rollouts effectively prevents the error accumulation that plagues baseline methods, enabling the generation of coherent and physically plausible manipulation sequences.
4.3 World Model for Policy Learning
Figure 6: Illustration of downstream policy learning with MIND-V.
To validate the utility of MIND-V as a training ground for embodied agents, we conduct a “Generation-to-Policy” experiment in the MimicGen simulation benchmark [mimicgen] (Fig. 6). We select three fine-grained manipulation tasks (Coffee, StackThree, Square) with a limited budget of 128 demonstration trajectories per task.
Setup  We first fine-tune a VLA (OpenVLA-OFT [openvla]) via imitation learning (IL) with 300 expert trajectories on target manipulation tasks as base policy. For comparison, the IL training strategy continues training this OpenVLA-OFT model using end-effector pose supervision ($\mathcal{L}_{\text{pose}}$) until convergence. In our approach, in addition to the standard pose loss, we leverage MIND-V to generate successful visual rollouts ($o$) conditioned on the task instructions, which serve as visual goals to guide the VLA policy. Specifically, the OpenVLA-OFT predicts an action sequence to execute a rollout ($\hat{o}$) in simulation engine, which is supervised by minimizing the visual discrepancy against MIND-V’s prediction:
$\mathcal{L}=\lambda_{\text{pix}}\|\hat{o}-o\|_{1}+\lambda_{\text{perc}}d_{\text{LPIPS}}(\hat{o},o)+\lambda_{\text{IL}}\mathcal{L}_{\text{pose}},$ (8)
where $\|\hat{o}-o\|_{1}$ and $d_{\text{LPIPS}}(\hat{o},o)$ denote the $L_{1}$ pixel reconstruction loss and LPIPS perceptual loss [LPIPS], respectively.
Table 3: Task success rates (%) under different training paradigms in the MimicGen simulation benchmark. Best results are bolded.
Methods Coffee StackThree Square Mean
Base policy 32.6 30.2 20.2 27.7
IL 37.4 36.7 26.1 33.4
MIND-V + IL 51.7 48.3 30.4 43.5
Dissusion  As detailed in Table 3, incorporating MIND-V significantly raises the performance ceiling of the VLA policy: the mean success rate improves from 27.7% (Base) to 43.5% with MIND-V guidance, far exceeding the 33.4% of IL continuous learning strategy. This result underscores a fundamental advantage of pixel-space world models: unlike latent-space IL or RL approaches that operate in abstract representations detached from visual reality, MIND-V generates futures in the same pixel space as the VLA’s pre-training data. This alignment allows the VLA to directly leverage its rich visual priors to learn from the world model’s imagined rollouts, effectively bridging the gap between generative world simulation and actionable policy learning [WMPO].